### v1

@@ Abstract
We introduce GraviBERT, a novel deep learning framework for inference of gravitational wave time series, which relies on an Inception-inspired multi-scale convolutional feature extractor with a transformer encoder along a suitable regression head. GraviBERT is trained in two stages: a BERT-style pretraining phase, where the model learns to predict masked segments in feature space in order to capture universal patterns and physics, followed by supervised fine tuning for accurate parameter estimation. This approach demonstrates impressive improvement across multiple metrics compared to training from scratch: On in-domain data, it reduces the mean absolute error for point-estimate parameter inference by up to 30%, and training convergence accelerates by up to $6\times$. Moreover, the mean relative precision of the inferred masses and distances at low SNR reaches the few-percent level, while the mean absolute error in the effective spin about $10^{-3}$. For domain adaptation to new detector noise profiles, the pretrained model demonstrates remarkable efficiency, converging up to $15\times$ faster on small target datasets and reduces estimation errors by up to $\sim 45$%, indicating that it learns sufficient detector-agnostic representations. Cross-approximant transfer demonstrates comparable transferability, achieving up to $44\%$ MAE reductions across all parameters and up to $15\times$ training speedups, with $R^{2}$ scores consistently exceeding $0.9$ for mass parameters at SNR = 10 compared to $0.74$ - $0.87$ when training from scratch. Notably, GraviBERT works directly with the noisy waveforms. The final regression head of the model can be adapted for various downstream tasks after pretraining, positioning it as a step towards foundation-style models in gravitational wave and multi-messenger astronomy.
@@ 1 Introduction
Time series analysis lies at the heart of numerous scientific and applied domains, from fundamental physics to finance and climate modeling. Traditionally, architectures such as autoregressive models, convolutional neural networks (CNNs), and recurrent neural networks (RNNs) have been widely employed for this task, with RNNs and their variants such as Long Short-Term Memory (LSTM) networks and Gated Recurrent Units (GRUs) the dominant choice due to their ability to model temporal dependencies in sequential data [Hochreiter1997, Cho2014].
@@
More recently, transformer architectures, originally developed for natural language processing, have redefined the landscape of time series modeling by offering significantly enhanced scalability and parallelism, as well as the ability to capture subtle global dependencies through the so–called self-attention mechanism [vaswani2017attention]. Yet, despite their growing success, the application of transformers to Bayesian parameter inference, particularly in the computationally demanding domain of gravitational wave astronomy, remains largely unexplored. In this work, we introduce a transformer-based architecture tailored for analysis of gravitational waveforms. Our approach bridges a gap in the field, demonstrating that transformers can excel in both efficiency and accuracy in this context, while it represents a step towards the development of foundation models for gravitational wave science which could be used for multi-modal tasks.
@@
The discovery of gravitational waves (GWs) marked a transformative moment in physics, opening an unprecedented observational window into the Universe [Abbott2016]. GWs are ripples of spacetime produced by the acceleration of compact astrophysical objects such as black holes and neutron stars, and are physically recorded by the detector as time-domain, oscillatory signals characterized modulated by the underlying dynamics of the source system. Beyond confirming General Relativity (GR) in the previously untested strong-field regime, GWs enable inference of key physical properties of their sources, such as component masses, spins, orbital dynamics, and geometrical configurations [TheLIGOScientificCollaboration2021]. This makes them invaluable for probing the nature of gravity over astrophysical and cosmological distances, the physics of stellar populations, testing theories related to dark matter or environmental effects, the physics of dark energy, or exotic fields [Berti2015, Saltas:2018fwy, Abac:2025saz, LISA:2022kgy, Barausse:2020rsu]. Reconstructing this wealth of information from noisy observational data constitutes a formidable inverse problem. Aside the challenges arising from the non-linear nature of Einstein’s equations, the low signal-to-noise ratios (SNR) and various instrumental and modelling systematics present hurdles with the waveform reconstruction. Recent work has shown that the computational challenges for higher-order modes may be somewhat mitigated through efficient search strategies [Estelles2024higher, Heinzel2024higher], though the problem remains significant for precessing systems with eccentricity. Consequently, the development of fast, accurate, and robust methods for GW signal analysis is essential to fully exploit the scientific potential of current and next-generation detectors.
@@
Machine learning (ML) has emerged as a promising tool to tackle the above challenge. GW data analysis proceeds in two stages: signal detection via matched filtering [abbott2016observation, allen2005chi2], which identifies candidate events by cross-correlating data with waveform templates, followed by parameter estimation using Bayesian inference [skilling2006nested, veitch2015parameter, ashton2019bilby] for confirmed detections. ML-based approaches can substantially accelerate GW data analysis [^1] We should note here that it is impossible to provide here an exhaustive list of previous works, and for reviews we refer to [cuoco2021enhancing, stergioulas2024machine, torres2024applications, ormiston2024machine, alyasiri2023ai, george2018deep, george2018deep2_detection, aasi2024contribution, huerta2022inference, krastev2020real, zhao2023space].. GW parameter estimation conventionally relies on matched filtering [abbott2016observation, wainstein1962extraction, allen2005chi2] and Bayesian inference [skilling2006nested, veitch2010bayesian, vandersluys2008parameter, cornish2015bayeswave, veitch2015parameter, ashton2019bilby, speri2024probing, cole2025sequential] (see also [thrane2019introduction, meyer2022computational] for reviews), requiring up to weeks per event. Recent neural network approaches achieve dramatic speedups: normalising flows accelerate posterior sampling [williams2021nested, williams2023importance, green2020gravitational, hu2025long], neural posterior estimation enables real-time inference through frameworks like DINGO, which solves the full 15-parameter BBH problem [dax2023dingo, dax2023nature], and related NPE methods [dax2021real, dax2023neural], conditional variational autoencoders reduce inference time from weeks to seconds [gabbard2022bayesian], and neural network emulators enable rapid likelihood evaluation [chua2019reducing, williams2021gaussian]. Recent transformer-based approaches for GWs include work on overlapping signal parameter estimation in Einstein Telescope (ET) scenarios [papalini2025transformer] and pretrained audio transformers [chatterjee2024transformer] (see also [Wang:2025aqk, Chan:2025pdf, Inguglia:2025cig] for similar applications with alternative architectures). Our work differs in its focus on unsupervised masked pretraining directly on noisy GW data for learning generalizable representations and demonstrating transfer learning across detector configurations, rather than addressing a specific astrophysical application. Score-based models [langendorff2023score] and density estimation methods using masked auto-regressive flows [green2021complete] offer additional approaches for robust Bayesian inference. Matched filtering itself presents substantial computational challenges, with template banks that scale prohibitively for higher-order modes and eccentricity in third-generation detectors [owen1999matched], making efficient alternatives critically urgent. However, ML methods typically depend on assumptions about noise, waveforms, and detector modeling, requiring retraining when configurations change. Most methods lack generalizability across signal types or parameter ranges, and their black-box nature can obscure uncertainty quantification and systematic errors.
@@
ML methods come with limitations – they usually depend on assumptions about the underlying noise, waveform or detector modelling. Training must be typically repeated whenever detector configurations, noise characteristics, or theoretical models evolve or change. In this regard, most methods lack a fair amount of generalizability across different signal types or parameter ranges, creating brittleness in operational deployment. Often, the black-box nature of these approaches obscures uncertainty quantification and makes systematic error detection challenging.
Transformers represent a revolutionary architecture that fundamentally transformed natural language processing [vaswani2017attention]. Their distinctive self-attention mechanism enables dynamic weighting of different temporal components within sequences while processing data in parallel, facilitating both computational scalability and the capture of long-range dependencies that traditional recurrent architectures struggle to model effectively. The architecture produces rich encoded representations (embeddings) that serve as versatile foundations for diverse downstream tasks through appropriate task-specific heads. This capability to generate transferable representations lies at the core of foundation models and enables multi-modal applications across domains. Pioneering examples of such transformer-based foundation models include BERT [Devlin2018] for bidirectional language understanding and the GPT series [Radford2018, Radford2019] for autoregressive generation, both of which have demonstrated remarkable versatility and performance across numerous tasks through pretraining and fine-tuning paradigms.
While full Bayesian posterior estimation remains the gold standard for final scientific analysis [veitch2015parameter, ashton2019bilby], our work focuses on demonstrating for the first time the effectiveness of transformer-based architectures with unsupervised pretraining for GW parameter inference. We use point estimates with MC Dropout uncertainty quantification [gal2016dropoutbayesianapproximationrepresenting] as a computationally efficient framework to evaluate our architectural and training novelties. Extension to full Bayesian inference, for example, by replacing the regression head with normalising flows—represents a natural but separate direction for future work. Our primary contributions therefore are: (i) demonstrating that transformer-based architectures effectively capture GW features, (ii) showing that unsupervised pretraining significantly improves training efficiency and accuracy, and (iii) establishing that pretrained models present with promising transfer learning capabilities. The name "GraviBERT" reflects our model’s conceptual inspiration from BERT’s masked pretraining approach, though adapted for continuous-valued time series rather than discrete tokens.
@@
We structure the paper as follows: In Section 2 we discuss our choice of waveform modelling and data generation. In Section 3 we introduce our model architecture, and proceed discussing its training in Section 4. In Section 5 we present our results and discuss our conclusions in Section 6.
@@ 2 Modelling and data generation
The GW signal from a binary merger, such as that of two black holes or neutron stars, exhibits three distinct evolutionary phases: inspiral, merger, and ringdown [Maggiore2008, Blanchet2014]. During the inspiral phase, the two compact objects orbit each other at relatively large separations, gradually losing orbital energy through gravitational radiation emission. As the objects approach closer separations, their gravitational interaction becomes highly nonlinear, marking the transition to the merger phase. Finally, the merger culminates when the two bodies coalesce into a single object, producing the characteristic peak amplitude in the observed waveform, followed by the ringdown phase where the remnant object settles to equilibrium [Pretorius2005, Campanelli2006]. The physical characteristics of these waveforms are essential for our subsequent analysis. At leading order in the post-Newtonian expansion, the GW strain can be approximated as [Blanchet2014]:
@@
where $\mathcal{M}_{c}\equiv(m_{1}m_{2})^{3/5}/(m_{1}+m_{2})^{1/5}$ is the chirp mass, $\omega(t)$ represents the center-of-mass orbital frequency, $\Phi(t)$ denotes the orbital phase, and $d$ is the luminosity distance to the source. The orbital phase evolution is governed by [Blanchet2014, Buonanno2009]:
@@
where $t_{c}$ denotes the coalescence time, $v\equiv\left(G(m_{1}+m_{2})\omega\right)^{1/3}$ is the orbital velocity, and $\chi_{\rm{eff}}\equiv(m_{1}s_{1}+m_{2}s_{2})/(m_{1}+m_{2})$ represents the effective spin parameter.
This formulation reveals several critical features for parameter estimation: (i) The luminosity distance provides an overall amplitude suppression scaling as $\sim 1/d$; (ii) At leading post-Newtonian order, the phase evolution is primarily governed by the binary’s masses; (iii) Higher-order corrections introduce mass-spin coupling terms that create parameter degeneracies during inference [Cutler1994, Vallisneri2008]. These degeneracies, particularly between mass and spin parameters, pose significant challenges for neural network-based parameter estimation and must be carefully addressed in our architecture design.
We utilise the Effective One Body (EOB) formalism through the SEOBNRv4 approximant [Bohe2017], which delivers accurate modeling of both inspiral and merger phases through a combination of analytical post-Newtonian theory and numerical relativity calibration. We will focus on binary black hole (BBH) mergers with mass ratios $q=m_{2}/m_{1}$ ranging between $q\sim 0.1-1$. These systems fall within the expected validity of our approximant and lie within the detectable range of current and next-generation GW observatories, such as Advanced LIGO and the ET.
GWs manifest as time-series strain measurements in interferometric detectors. Given the limited number of confirmed detections from current observatories [Abbott2021], training robust neural networks necessitates extensive simulated datasets. For waveform generation, we employ the open-source PyCBC library [Nitz2024], which provides access to state-of-the-art waveform approximants. Our training dataset is constructed using a grid-based approach over source masses, spins and luminosity distance. We exclude source orientation or other parameters from our grid, as we are here interested in a first proof-of-concept analysis. We employ a fixed sampling rate of 4096 Hz with a lower frequency cutoff of $f_{\rm low}=20$ Hz and simulate ET detector noise using as our reference the ET-B power spectral density (PSD) [Hild2011]. While ET typically uses $f_{\rm low}=5$ Hz to maximize sensitivity, we use 20 Hz for this proof-of-concept to reduce computational costs and waveform duration, noting that the lower cutoff primarily affects low-mass systems and does not fundamentally limit the applicability of our method.
@@
The distance $d$ denotes an effective luminosity distance used to rescale the signal amplitude and enforce a fixed target SNR. When detector noise is included, $d$ is adjusted for each waveform realization using the scaling $h(t)\propto 1/d$. We focus on two fixed SNR values (10 and 30) to enable controlled comparison of model performance under different (conservative) noise conditions. This assumption must be compared with the continuous SNR distributions expected in realistic astrophysical observations. Extension to variable SNR distributions motivated by astrophysical populations is a direction for future work.
@@
We emphasise that we focus on the four intrinsic parameters and do not model extrinsic parameters such as sky location ($\alpha,\delta$), inclination angle ($\iota$), polarization angle ($\psi$), or coalescence phase ($\phi_{c}$). Consequently, we do not apply the detector antenna pattern function $F^{+}(\alpha,\delta,\psi,t)$ and $F^{\times}(\alpha,\delta,\psi,t)$ that would project the GW polarizations onto the detector frame. Our strain data represents the waveform $h(t)$ in the source frame at optimal orientation, rather than the detector-frame strain $s(t)=F^{+}h_{+}(t)+F^{\times}h_{\times}(t)$ that would be observed in a real interferometer. This simplification significantly reduces the complexity of the parameter space but limits direct applicability to observational data. However, it serves our primary objective: demonstrating that unsupervised pretraining and transfer learning are effective strategies for GW analysis. The objective of this work are orthogonal to whether the full parameter space is modelled, that is, the same pre-training approach can be applied to architectures handling all 15 parameters.
We note that the $(m_{1},m_{2})$ parameterization used here exhibits stronger correlations than alternative parameterizations such as chirp mass and mass ratio. Future implementations could benefit from using $({\cal M}_{c},q)$ or similar parameterizations that align more naturally with the physical variations encoded in the waveform, potentially improving training efficiency and reducing parameter degeneracies.
We generate training data on a regular grid over the parameter space rather than drawing randomly from astrophysically motivated priors. While production systems typically use prior-based sampling, grid-based generation offers advantages for our proof-of-concept study: (i) it ensures uniform coverage of the parameter space, (ii) it enables systematic evaluation of model performance across parameter ranges, and (iii) it provides a controlled setting for isolating the effects of our pretraining and transfer learning strategies. We emphasise that the network performs regression on continuous values, i.e the grid structure in training data does not reduce this to a classification problem, as the network must interpolate between grid points and is trained with mean squared error loss on continuous targets. This represents a simplification compared to realistic inference scenarios.

### v2

@@ Abstract
We introduce GraviBERT, a novel deep learning framework for gravitational wave inference, built on a multi-scale feature extractor with a transformer encoder and a suitable regression head. A key novelty of GraviBERT is its staged training: a BERT-style self-supervised pretraining phase to learn transferable representations, followed by supervised fine-tuning on labeled data. GraviBERT demonstrates consistent transfer learning across detector configurations and waveform models. On in-domain data, pretraining reduces the MAE by up to 31% and accelerates convergence by $\sim 6.6\times$, with mean relative precision for point estimates reaching the few-percent level and MAE in effective spin of $\sim 10^{-3}$ at SNR = 10. For domain adaptation to new detector noise profiles, the pretrained model converges up to $15\times$ faster on small target datasets and reduces estimation errors by up to $\sim 47$%, demonstrating detector-agnostic learning. Cross-waveform approximant transfer achieves up to $44\%$ MAE reductions and up to $15\times$ training speedups, with $R^{2}$ scores consistently exceeding $0.9$ for mass parameters at SNR = 10 compared to $0.74$ - $0.87$ when training from scratch. GraviBERT works directly with noisy waveforms, and in its current form quantifies predictive uncertainty through MC dropouts. After pretraining, the regression head could be adapted to multiple downstream inference tasks in gravitational-wave astronomy.
@@ 1 Introduction
Time series analysis lies at the heart of numerous scientific and applied domains, from fundamental physics to finance and climate modelling. Traditionally, auto-regressive models and recurrent neural networks (RNNs) have been widely employed for this task, with RNNs and their variants such as Long Short-Term Memory (LSTM) networks [30] the dominant choice.
@@
More recently, transformer architectures, originally developed for natural language processing, have redefined the landscape of time series modelling by offering significantly enhanced scalability and parallelism, as well as the ability to capture subtle global dependencies through the so–called self-attention mechanism [48]. Yet, despite their growing success, the application of transformers to the computationally demanding domain of gravitational-wave astronomy, remains mostly unexplored. In this work, we introduce a transformer-based architecture tailored for analysis of gravitational waveforms. Our approach demonstrates that an architecture based on a transformer encoder combined with self-supervised pretraining can improve both the efficiency and accuracy for gravitational-wave parameter estimation. This approach suggests a potential pathway towards foundation-style models for gravitational-wave science that could be extended to multi-modal tasks.
@@
The discovery of gravitational waves (GWs) marked a key moment in physics, opening an unprecedented observational window into the Universe [1, 4]. This makes them invaluable for probing the nature of gravity over astrophysical and cosmological distances, the physics of stellar populations, testing theories related to dark matter or environmental effects, the physics of dark energy, or exotic fields [9, 40, 6, 8]. Reconstructing this wealth of information from noisy observational data constitutes a challenging inverse problem. Aside the challenges arising from the non-linear nature of Einstein’s equations, the low signal-to-noise ratios (SNR) and various instrumental and modelling systematics present hurdles with the waveform reconstruction. Consequently, the development of fast, accurate, and robust methods for GW signal analysis is essential to fully exploit the scientific potential of current and next-generation detectors.
@@
Machine learning (ML) has emerged as a promising tool to tackle the above challenge. GW data analysis relies on parameter estimation using Bayesian inference [42, 49, 7] for confirmed detections. ML-based approaches can substantially accelerate GW data analysis [^1] We should note here that it is impossible to provide here an exhaustive list of previous works, and for reviews we refer to [18, 43].. GW parameter estimation conventionally relies on matched filtering [1, 5] and Bayesian inference [42, 50, 47, 17, 49, 7, 10] (see also [45] for reviews), requiring up to days per event. Recent neural network approaches achieve dramatic speedups: normalising flows accelerate posterior sampling [52, 53, 28], neural posterior estimation enables real-time inference through frameworks like DINGO, which solves the full 15-parameter BBH problem [22, 21], and related NPE methods [22, 20], conditional variational autoencoders reduce inference time from weeks to seconds [25], and neural network emulators enable rapid likelihood evaluation [16, 52]. Recent transformer-based approaches for GWs include work on overlapping signal parameter estimation in Einstein Telescope (ET) scenarios [37] and pretrained audio transformers [15] (see also [51, 14, 31] for similar applications with alternative architectures). Our work differs in its focus on self-supervised masked pretraining directly on noisy GW data for learning generalisable representations and demonstrating transfer learning across detector configurations, rather than addressing a specific astrophysical application. Very recently, Ref. [34] introduced Dingo-T1, a transformer-encoder–based model that achieves flexibility via tokenisation of frequency-segmented data, enabling adaptation at inference time to different detector configurations. In contrast, GraviBERT combines a multi-scale feature extractor with a transformer encoder and advances a distinct transfer-learning paradigm: BERT-style masked pretraining on unlabeled GW data to learn transferable representations of the underlying physics. These representations can, in principle, support multiple downstream tasks and accelerate adaptation to new detectors and waveform approximants, demonstrating for the first time that self-supervised pretraining enables efficient transfer learning in GW analysis.
@@
ML methods come with limitations – they usually depend on assumptions about the underlying noise, waveform or detector modelling. Training must be typically repeated whenever detector configurations, noise characteristics, or theoretical models evolve or change. In this regard, most methods lack a fair amount of generalisability across different signal types or parameter ranges, creating brittleness in operational deployment. Often, the black-box nature of these approaches obscures uncertainty quantification and makes systematic error detection challenging.
Transformers revolutionised natural language processing through their self-attention mechanism, which enables dynamic weighting of temporal components while processing data in parallel [48]. This architecture produces rich embeddings that serve as versatile foundations for diverse downstream tasks, forming the basis of foundation models like BERT [23] and the GPT series [38, 39], which achieve broad performance through pretraining and fine-tuning paradigms.
Building on this paradigm, we develop GraviBERT, which adapts transformer architectures to gravitational wave analysis. Our multi-scale convolutional feature extractor downsamples the time-domain waveform while progressively upsampling feature dimensions, hence learning local temporal features. The compressed representations are passed to the transformer encoder, which encodes transferable GW features through self-attention. Our primary contributions are: (i) demonstrating that this architecture captures transferable representations of GW signals, (ii) showing that self-supervised pretraining improves training efficiency and accuracy compared to training from scratch, and (iii) establishing promising transfer learning capabilities for domain and waveform adaptation. The name GraviBERT reflects our model’s conceptual inspiration from BERT’s masked pretraining approach, adapted for continuous-valued time series.
Our focus is on demonstrating for the first time the effectiveness of transformer-based architectures with self-supervised pretraining for GW parameter inference. Therefore, we work in the simplified setting of point estimates backed by with MC Dropout uncertainty quantification [26] as a computationally efficient framework to evaluate our architectural and training novelties. Extension to full Bayesian inference represents a natural, but separate direction which deserves its own study.
@@
We structure the paper as follows [^2] Our code and pretrained models are publicly available at https://zenodo.org/records/18671949.: In Section 2 we discuss our choice of waveform modelling and data generation. In Section 3 we introduce our model architecture, and proceed discussing its training in Section 4. In Section 5 we present our results and discuss our conclusions in Section 6.
@@ 2 Modelling and data generation
The GW signal from a binary merger, such as that of two black holes or neutron stars, exhibits three distinct evolutionary phases: inspiral, merger, and ringdown [36, 11]. At leading order in the post-Newtonian expansion, the GW strain can be approximated as [11],
@@
where $\mathcal{M}_{c}\equiv(m_{1}m_{2})^{3/5}/(m_{1}+m_{2})^{1/5}$ is the chirp mass, $\omega(t)$ represents the centre-of-mass orbital frequency, $\Phi(t)$ denotes the orbital phase, and $d$ is the luminosity distance to the source. The orbital phase evolution is governed by [11, 13]:
@@
where $t_{c}$ denotes the coalescence time, $v\equiv\left(G(m_{1}+m_{2})\omega\right)^{1/3}$ is the orbital velocity, and $\chi_{\rm{eff}}\equiv(m_{1}s_{1}+m_{2}s_{2})/(m_{1}+m_{2})$ represents the effective spin parameter. Notice that higher-order corrections introduce mass-spin coupling terms that create parameter degeneracies during inference [19, 46].
We utilise the Effective One Body (EOB) formalism through the SEOBNRv4 approximant [12], which delivers accurate modelling of both inspiral and merger phases through a combination of analytical post-Newtonian theory and numerical relativity calibration. We will focus on binary black hole (BBH) mergers with mass ratios $q=m_{2}/m_{1}$ ranging between $q\sim 0.1-1$. These systems fall within the expected validity of our approximant and lie within the detectable range of current and next-generation GW observatories, such as Advanced LIGO and the ET.
GWs manifest as time-series strain measurements in interferometric detectors. Given the limited number of confirmed detections from current observatories [2], training robust neural networks necessitates extensive simulated datasets. For waveform generation, we employ the open-source PyCBC library, which provides access to state-of-the-art waveform approximants. Our training dataset is constructed using a grid-based approach over source masses, spins and luminosity distance. We exclude source orientation or other parameters from our grid, as we are here interested in a first proof-of-concept analysis. We employ a fixed sampling rate of 4096 Hz with a lower frequency cutoff of $f_{\rm low}=20$ Hz and simulate ET detector noise using as our reference the ET-B power spectral density (PSD) [29]. While ET typically uses $f_{\rm low}=5$ Hz to maximise sensitivity, we use 20 Hz for this proof-of-concept to reduce computational costs and waveform duration, noting that the lower cutoff primarily affects low-mass systems and does not fundamentally limit the applicability of our method.
@@
The distance $d$ denotes an effective luminosity distance used to rescale the signal amplitude and enforce a fixed target SNR. When detector noise is included, $d$ is adjusted for each waveform realization using the scaling $h(t)\propto 1/d$. We focus on two fixed SNR values (10 and 30) to enable controlled comparison of model performance under different (conservative) noise conditions. This assumption must be compared with the continuous SNR distributions expected in realistic astrophysical observations. We fix the length of our input data to $4096$ using appropriate padding as needed.
@@
We emphasise that we focus on the four intrinsic parameters and do not model extrinsic parameters such as sky location ($\alpha,\delta$), inclination angle ($\iota$), polarization angle ($\psi$), or coalescence phase ($\phi_{c}$). Consequently, we do not apply the detector antenna pattern function $F^{+}(\alpha,\delta,\psi,t)$ and $F^{\times}(\alpha,\delta,\psi,t)$ that would project the GW polarizations onto the detector frame. Our strain data represents the waveform $h(t)$ in the source frame at optimal orientation, rather than the detector-frame strain $s(t)=F^{+}h_{+}(t)+F^{\times}h_{\times}(t)$ that would be observed in a real interferometer. This simplification significantly reduces the complexity of the parameter space but limits direct applicability to observational data. However, it serves our primary objective: demonstrating that self-supervised pretraining and transfer learning are effective strategies for GW analysis. The objectives of this work are orthogonal to whether the full parameter space is modelled, that is, the same pretraining approach can be applied to architectures handling all 15 parameters.
We note that the $(m_{1},m_{2})$ parameterisation used here exhibits stronger correlations than alternative parameterisations such as chirp mass and mass ratio. Future implementations could benefit from using $({\cal M}_{c},q)$ or similar parameterisations that align more naturally with the physical variations encoded in the waveform, potentially improving training efficiency and reducing parameter degeneracies.
We generate training data on a regular grid over the parameter space rather than drawing randomly from astrophysically motivated priors. While production systems typically use prior-based sampling, grid-based generation offers advantages for our proof-of-concept study: (i) it ensures uniform coverage of the parameter space, (ii) it enables systematic evaluation of model performance across parameter ranges, and (iii) it provides a controlled setting for isolating the effects of our pretraining and transfer learning strategies. The network performs regression on continuous values, i.e the model outputs continuous predictions, and is not constrained to return grid values. However, since our test set shares the same grid structure, we do not explicitly evaluate off-grid generalisation.